OpenStudy (rahulmr):

Given \[\int\limits_{0}^{4}f(x)dx=8\]. Find \[\int\limits_{0}^{12}(2f(3x)-x)dx\]. I don't know how should i do this question. Can someone help me out. Thanks.

OpenStudy (kainui):

What have you tried? Play around with it with some algebra a little bit and show me what you get.

OpenStudy (rahulmr):

\[2\int\limits_{0}^{12}f(3x)dx-\int\limits_{0}^{12}(x)dx\] I don't know what to do next.

OpenStudy (loser66):

the second one is ok, right? Now the first one you need find \[2\int_0^{12} f(3x) dx\] Let u =3x, du /3 =dx x= 0 , u =0 x=12 u =4, then you have \ \[2\int_0^4 f(u) du =2*8\] Don't forget + the second part

OpenStudy (rahulmr):

Thanks, But how can i get rid of the 3 when i'll sub the 3x back into the equation

jimthompson5910 (jim_thompson5910):

if x = 12, then u = 3x = 3*12 = 36

OpenStudy (loser66):

oh yeah!! my bad!! I am so sorry Thank you so much @Mr. Jim :)

OpenStudy (loser66):

I should go to bed now. :)

OpenStudy (rahulmr):

Okay. Thanks for the help

jimthompson5910 (jim_thompson5910):

Are you sure the 3x is there? If it was x/3, then it would work. I'm thinking there's a typo somewhere.